OpenStudy (anonymous):

The u.s senate has 100 members. After a certain election there were four more democrats than republicans with no other parties represented. How many members of each party were their in the senate?

jigglypuff314 (jigglypuff314):

x -> dem y -> rep x + y = 100 y = x + 4

OpenStudy (anonymous):

ok so what is the best way to figure this out?

jigglypuff314 (jigglypuff314):

substitution. plug in x+4 for y into x+y =100 to get x+x+4 = 100 solve for x

OpenStudy (anonymous):

so what numbers do i plug in? is it what equals 100?

jigglypuff314 (jigglypuff314):

I plugged in y for you just solve for x now :) x+x+4 = 100 2x + 4 = 100

OpenStudy (anonymous):

48? so where do iput 48?

jigglypuff314 (jigglypuff314):

x = democrats = 48

OpenStudy (anonymous):

ok how do i find the republicans?

jigglypuff314 (jigglypuff314):

then plug 48 for x back into y = x + 4 and solve for y :)

OpenStudy (anonymous):

52?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

it says its wrong?

jigglypuff314 (jigglypuff314):

oh crap sorry, I mixed up the x and ys it's the right numbers,but democrats would have the larger number

OpenStudy (anonymous):

thanks :)